OpenStudy (anonymous):

how do you use the substitution method for 5x+3y=7 and y=x-3

OpenStudy (anonymous):

You know y=x+3 so put x-3 in for y in the other equations and solve for x 5x+3(x-3)=7 then use x to solve for y

OpenStudy (anonymous):

x-3*****

OpenStudy (anonymous):

ok, so after you write 5x+3(x-3)=7 then what do you do from there?

OpenStudy (lilg132):

you replace y with the equation given so 5x + 3y =7 y = x - 3 so it will now be 5x + 3(x - 3) = 7 expand it 5x + 3x - 9 = 7 bring the numbers to one side and letters on the other 5x + 3x = 7 + 9 8x = 16 now divide 8x by 8 to get x on its own but you have to do it to both sides 8x / 8 = 16 / 8 x = 2

OpenStudy (lilg132):

now we know what the value of x is we can work out y 5x + 3y =7 so 5(2) + 3y = 7 expand the bracket 10 + 3y = 7 now move numbers to one side and letters to the other 3y = 7 - 10 3y = -3 divide both sides by 3 to get y on its own 3y / 3 = -3 / 3 y = -1 DONT FORGET TO CLICK BEST RESPONSE =)

OpenStudy (anonymous):

ok thanks so much! ok so i have a couple more i might need help with....so how bout 3x+2y=11 and x+y=4?

OpenStudy (anonymous):

can someone please help!